OpenStudy (anonymous):

Write an expression in simplest form for the volume of a cylinder with a radius of 4a2 and a height of 5a. Use 3.14 for pi. A. 251.2a5 B. 62.8a3 C. 240a4

OpenStudy (hoblos):

V= pi*h*r^2 = 3.14*5a*16a^4 = 251.2a5

OpenStudy (hoblos):

A